[Brahms-l] Shift report 20060614 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Wed Jun 14 2006 - 00:09:34 EDT
Shifter leader was dp
Shifters were te


We had beam since 6:30 PM. We took data for 4 hrs and 15 minutes. MCR did polarization measurement several times and we needed to stop taking data during that period. Anyway, besides that there were smooth beam..
_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Jun 14 00:10:43 2006

This archive was generated by hypermail 2.1.8 : Wed Jun 14 2006 - 00:10:58 EDT